OpenStudy (anonymous):

Given the equation 3x - 4y = 8, which equation below would cause a consistent-dependent system? 3x + 4y = -8 6x - 8y = 12 9x - 12y = 24 16x + 12y = -10

OpenStudy (anonymous):

@linh412986 can u give me the answer?

OpenStudy (anonymous):

is it C?

OpenStudy (anonymous):

OK. let's see what is a consistent-dependent system's requirement? Consistent: at least one solution Dependent: infinite number of solutions So consistent-dependent means that linear equation systems must have infinite number of solutions. You can use many methods to check this requirement. For me, this case will be solved faster when using definition of coincides line in 2D coordination. We have: 3x - 4y = 8 \[y = \frac{3}{4}x-2\] In order for other lines coincides to y = (3/4) * x - 2, we must have the slope of them is equal to 3/4 The rest job is to check the slope of 4 options. After checking, I found that: C is the only option. In fact, you can see C is just 3x - 4y = 8 and then scaled by factor 3.

OpenStudy (anonymous):

so its C?

OpenStudy (anonymous):

As I said in my answer, do not ask me again like that ^^

OpenStudy (anonymous):

ok thx